An environmental chemist needs a carbonate buffer

of pH 10.00 to study the effects of acid rain on

limestone-rich soils. How many grams of Na2CO3

must she add to 1.5 L of freshly prepared 0.20 M

NaHCO3

to make the buffer? pKa of HCO3–

is 10.


1
Expert's answer
2021-06-02T06:31:02-0400

pH= 10.0

[NaHCO3] = 0.20M

Ka= 4.7 x 10-11

pKa= -log Ka

pKa= -log (4.7 x 10-11) = 10.33


Using Henderson Hasselbalch equation:

pH= pKa + log[Na2CO3]/[NaHCO3]

10.0 = 10.33 + log [Na2CO3]/0.20


Log[Na2CO3]/0.20 = 10.0 - 10.33 = -0.33

[Na2CO3]/0.20 = 10-0.33

[Na2CO3] = 0.20 x 0.47 = 0.094M


Moles of Na2CO3= C x V

= 0.094M x 1.5L

= 0.141mole


Molar mass of Na2CO3= 106gmol-1

Mass= mole x molar mass

= 0.141 x 106 = 14.95g